OpenStudy (anonymous):

f(x) and g(x) has the same gradient at point (p,q). f(x) is a normal to the curve h(x). Now the question: Does f(x) normal to the curve h(x) at (p,q)?

OpenStudy (anonymous):

Original question from here, question 1 - part ii) http://dl.dropbox.com/u/63664351/STEP%201%20Final%202011.pdf Following is the solution: http://dl.dropbox.com/u/63664351/Final%20STEP%202011%20solutions.pdf Look at page 5 I quote "The gradient of the straight line is -a/b as before, so as this line is normal to the curve at thepoint (p, q)..."

OpenStudy (anonymous):

In here,\[f(x)=ax+by=1\]\[g(x)=\frac{a}{x}+\frac{b}{y}=1\]\[h(x)=\frac{a}{x}-\frac{b}{y}=1\]

OpenStudy (anonymous):

As you can read, in part (i) f(x) has the same gradient to g(x) at point (p,q). In the question on part (ii) the question only states that f(x) is a normal to h(x) but yet in the solution, I don't know how, they just write that they are normal at (p,q).

OpenStudy (experimentx):

-a/x^2 + b/y^2 dy/dx = 0 => dy/dx = ay^2/bx^2 solve: -a/b*dy/dx = -1 find the value of dy/dx from by putting vaues of x and y here http://www.wolframalpha.com/input/?i=solve+ax%2Bby%3D1%2C+a%2Fx-b%2Fy%3D1&dataset=

OpenStudy (experimentx):

yes ... that is the condition we imposed my doing this m1*m2 = -1

OpenStudy (experimentx):

it think you can do long and roundabout method by the process i mentioned above. but the solution given in your document is much shorter and better.

OpenStudy (anonymous):

I agree that \( a^2−b^2=1/2 \) But why f(x) normal with h(x) at \( (p,q) \) ? The very point where f(x) and g(x) has the same gradient!

OpenStudy (experimentx):

it it given that point p,q lies on the line as well as the curve a/x + b/y = 1

OpenStudy (anonymous):

Yes, but in part 2 the equation is a/x - b/y =1 There is no mention whatsoever about (p,q) on that curve.

OpenStudy (experimentx):

a/p = -b/q since p,q lies in the curve it must satify a/p - b/q = 1 => a/p - b/q = 1 => 2a/p = 1

OpenStudy (experimentx):

that was just simple substitution that looked like a/x +b/y = 1

OpenStudy (anonymous):

Okay lets step back for a second and confirm that we're on the same ground. - What does it means by normal to the curve?

OpenStudy (experimentx):

the product of slope must be -1

OpenStudy (anonymous):

Yes but what does it really means? Can you describe it graphically?

OpenStudy (experimentx):